# Video Frame Interpolation with Region-Distinguishable Priors from SAM

### v1

@@ Abstract
In existing Video Frame Interpolation (VFI) approaches, the motion estimation between neighboring frames plays a crucial role. However, the estimation accuracy in existing methods remains a challenge, primarily due to the inherent ambiguity in identifying corresponding areas in adjacent frames for interpolation. Therefore, enhancing accuracy by distinguishing different regions before motion estimation is of utmost importance. In this paper, we introduce a novel solution involving the utilization of open-world segmentation models, e.g., SAM (Segment Anything Model), to derive Region-Distinguishable Priors (RDPs) in different frames. These RDPs are represented as spatial-varying Gaussian mixtures, distinguishing an arbitrary number of areas with a unified modality. RDPs can be integrated into existing motion-based VFI methods to enhance features for motion estimation, facilitated by our designed play-and-plug Hierarchical Region-aware Feature Fusion Module (HRFFM). HRFFM incorporates RDP into various hierarchical stages of VFI’s encoder, using RDP-guided Feature Normalization (RDPFN) in a residual learning manner. With HRFFM and RDP, the features within VFI’s encoder exhibit similar representations for matched regions in neighboring frames, thus improving the synthesis of intermediate frames. Extensive experiments demonstrate that HRFFM consistently enhances VFI performance across various scenes.
@@
Figure 1: The first two columns: overlay inputs and the ground truth frame. Middle two columns: motion field (from first to second frame) by VFIformer [29] and corresponding interpolation. The last two columns: motion field and interpolated frame by enhancing VFIformer with our strategy using RDPs. Our strategy results in more satisfactory motion estimation, and thus better interpolation results.
1 Introduction
@@
Video frame interpolation (VFI) represents a classic low-level vision task with the objective of augmenting video frame rates by generating intermediary frames that do not exist between consecutive frames. This technique has a wide range of practical applications, such as novel view synthesis [10], video compression [28], and cartoon creation [45]. Nevertheless, frame interpolation continues to present unsolved challenges, including issues related to occlusions, substantial motion, and alterations in lighting conditions. Enhancing the performance of existing VFI frameworks in an efficient manner poses a significant challenge within both the research and industrial communities.
@@
The referenced VFI research can be broadly categorized into two main approaches: motion-free [35, 6, 2, 31] and motion-based [15, 25, 32, 1, 33, 39, 40, 29], depending on whether they incorporate motion cues like optical flow. Motion-free models typically utilize methods such as kernel prediction or spatial-temporal decoding, which are effective while have limitations, such as being restricted to interpolating frames at fixed time intervals, and their runtime scales linearly with the number of desired output frames. On the other end of the spectrum, motion-based approaches establish dense correspondences between frames and employ warping techniques to generate intermediate pixels. Due to the explicit modeling of temporal correlations, motion-based strategies are more flexible. Moreover, with recent advancements in optical flow technology [14, 13, 47, 48], motion-based interpolation’s accuracy has evolved into a promising framework.
@@
Motion estimation between adjacent frames is a pivotal aspect of motion-based Video Frame Interpolation (VFI). Nevertheless, achieving precise estimation accuracy in existing methods remains a formidable challenge, primarily due to the inherent ambiguity in identifying corresponding areas in adjacent frames for interpolation. This challenge becomes more pronounced when there is a substantial temporal gap in the target video. Previous research has predominantly focused on enhancing estimation accuracy by laboriously evolving network structures. In this paper, we posit that, in addition to network evolution, it is of paramount importance to enhance accuracy by differentiating between various regions prior to the motion estimation process.
In this paper, we present an innovative approach by introducing Region-Distinguishable Priors (RDPs) into motion-based VFI frameworks. These priors are derived from the existing open-source Segment-Anything Model (SAM) [20] with minimal impediments. Furthermore, we propose a new Hierarchical Region-aware Feature Fusion Module (HRFFM), which is designed to enhance the VFI framework’s encoder, as illustrated in Fig. 2, to refine the corresponding features used in motion estimation. The HRFFM is a plug-and-play module that seamlessly integrates with various motion-based VFI methods without introducing a significant increase in network parameters.
The formulation of RDP from SAM is not trivial, as RDP is required to differentiate objects with an arbitrary number, while the output of SAM lacks a countable property. To make optimal use of the segment outputs from SAM and provide them with the ability to distinguish multiple objects of the same dimensions, we have devised a novel Gaussian embedding strategy for the SAM outputs. We employ the Segment-Anything Model to produce instance segmentations for two input frames and utilize spatial-varying Gaussian mixtures to transform them into higher-dimensional RDPs. This representation has been demonstrated to outperform naive one-hot encoding or other learnable embedding alternatives.
The obtained RDP are integrated into the encoder of the target VFI model, with the primary goal of achieving regional consistency between neighboring frames in VFI. This means that the features of a specific region in two consecutive frames should be similar, which aids in the subsequent motion estimation process. To achieve this objective, HRFFM incorporates RDP into the target model’s hierarchical feature spaces and performs RDP-guided Feature Normalization (RDPFN) in a residual learning fashion to bring target features to desired states. RDPFN is novelly designed to simultaneously harness long- and short-range dependencies to fuse the RDP and image content, enabling the accurate estimation of regional normalization parameters.
Extensive experiments are conducted on public and well-recognized datasets and various VFI networks. It’s verified that our algorithm can bring stable performance improvement consistently on multiple datasets and models. Our strategy produces better motion modeling even with large motion scales, and thus enhances interpolated results (see Fig. 1). In summary, our contribution is three-fold.
  • We underscore the significance of distinguishing different regions within frames to enhance motion estimation and ultimately improve the performance of VFI. To achieve this, we have innovatively devised a novel formulation for RDP using a Gaussian embedding strategy based on the output of SAM.
  • A new Hierarchical Region-aware Feature Fusion Module is designed to incorporate RDPs into the target model’s encoder, and it is a general strategy for different networks.
@@
2 Related Work
2.1 Video Frame Interpolation
@@
The current VFI methods can be broadly categorized into two groups: motion-free and motion-based approaches. Motion-free methods typically create intermediate frames by directly concatenating input frames. Such methods can be further classified into two types: directly-generated methods [7, 11, 19, 27] and kernel-based methods [3, 4, 9, 22, 36, 37, 38, 43] concerning the generation of intermediate frames.Despite their simplicity, these methods lack a robust modeling of motion, making it challenging to align corresponding regions between intermediate frames and input frames. This limitation often results in image blur and the presence of artifacts [23].
@@
Motion-aware methods explicitly model motion, often represented by optical flow, between two frames to enhance the alignment of distinguishable region information from input frames to intermediate frames. Some early approaches focused solely on predicting inter-frame motion for pixel-level alignment [16, 26, 24]. Subsequent works [29, 32, 33, 39, 40, 49, 44, 42] have introduced separate modules for explicit motion modeling and motion refinement through synthesis, thereby enhancing overall performance. While the current state-of-the-art method has achieved impressive results, these systems still cannot handle practical challenges and need further performance improvement [51]. Our proposed method offers a novel perspective by incorporating Region Distinguishable Priors into motion-based VFI. Our designed play-and-plug Hierarchical Region-aware Feature Fusion Module provides a straightforward and efficient approach to improving VFI features via RDPs.
2.2 Segment Anything Model (SAM)
@@
The foundational Computer Vision (CV) model for Segment Anything, known as SAM [20], was recently unveiled. SAM is a substantial Vision Transformer (ViT)-based model that underwent training on an extensive visual corpus (SA-1B). Its capabilities in segmentation have shown promise across various scenarios, underscoring the significant potential of foundational models in the realm of CV. This development marks a groundbreaking stride toward achieving visual artificial general intelligence.
@@
SAM has demonstrated its versatility across a spectrum of CV tasks, extending its assistance beyond segmentation. Tasks such as image synthesis [50] and video super-resolution [30] have all benefited from SAM’s capabilities. In a pioneering effort, we’ve explored SAM’s potential in VFI, marking the first attempt to apply SAM to this domain. Extensive experiments substantiate that SAM significantly enhances the effectiveness of VFI.
3 Method
@@
In this section, we first provide the overview of our strategy in Sec. 3.1. Then, two vital components in our framework, i.e., the formulation of Region-Distinguishable Priors and the design of HRFFM, will be elaborated in Sec. 3.2 and 3.3, respectively. One significant component in HRFFM, i.e., RDP-guided Feature Normalization (RDPFN), will be introduced in Sec. 3.4.
3.1 Overview
@@
Implementation. Given $I_{0}$ and $I_{1}$, we first obtain their SAM outputs as $M_{0}$ and $M_{1}$. Then, $M_{0}$ and $M_{1}$ (${M_{0},M_{1}}\in\mathbb{R}^{H\times W\times 1}$) are transformed into the desired Region-Distinguishable Priors (RDPs) that can distinguish different regions in frames with a unified representation dimension. Thus, Eq. 1 can be written as
@@
3.2 Region-Distinguishable Priors(RDPs)
The drawback of SAM outputs for VFI. The original SAM model provides segmentation outputs for all instances within an image. SAM generates masks for frames, with each pixel value representing an object. Its remarkable segmentation capabilities make it a valuable choice as a region-distinguishable prior. Over time, several variants of SAM have been introduced, enhancing its capabilities, including semantic and panoptic segmentation when combined with other models. However, SAM’s output has limitations when it comes to representing objects with arbitrary numbers, a requirement for RDP. The semantic one-hot embedding is constrained by semantic categories, and the instance one-hot embedding assumes a maximum instance number, making it unable to accommodate new instances during real-world evaluation. Consequently, there is a need to transform SAM’s output to make it more suitable for RDPs.
Mixture Gaussian embedding strategy. We posit that the representations of segmentation priors can be conceptualized as distributed sampling results with distinct parameters across different regions of an image. These parameters enable the discrimination of regions and the alignment of the same region across multiple frames. In particular, each segmented area can be interpreted as a sampling result from a Gaussian distribution characterized by individual parameters. To facilitate this corresponding sampling process, we begin by establishing a codebook $\mathcal{C}$ that comprises a range of Gaussian parameters, encompassing both mean and variance. Subsequently, each object identified by the SAM output can retrieve its specific Gaussian parameters via a hashing mechanism. Therefore, the transformation procedure can be written as:

### v2

# Restoration-Oriented Video Frame Interpolation with Region-Distinguishable Priors from SAM

@@ Abstract
In existing restoration-oriented Video Frame Interpolation (VFI) approaches, the motion estimation between neighboring frames plays a crucial role. However, the estimation accuracy in existing methods remains a challenge, primarily due to the inherent ambiguity in identifying corresponding areas in adjacent frames for interpolation. Therefore, enhancing accuracy by distinguishing different regions before motion estimation is of utmost importance. In this paper, we introduce a novel solution involving the utilization of open-world segmentation models, e.g., SAM2 (Segment Anything Model2) for frames, to derive Region-Distinguishable Priors (RDPs) in different frames. These RDPs are represented as spatial-varying Gaussian mixtures, distinguishing an arbitrary number of areas with a unified modality. RDPs can be integrated into existing motion-based VFI methods to enhance features for motion estimation, facilitated by our designed play-and-plug Hierarchical Region-aware Feature Fusion Module (HRFFM). HRFFM incorporates RDP into various hierarchical stages of VFI’s encoder, using RDP-guided Feature Normalization (RDPFN) in a residual learning manner. With HRFFM and RDP, the features within VFI’s encoder exhibit similar representations for matched regions in neighboring frames, thus improving the synthesis of intermediate frames. Extensive experiments demonstrate that HRFFM consistently enhances VFI performance across various scenes.
I Introduction
@@
Figure 1: The first two columns: overlay inputs and the ground truth frame. Middle two columns: motion field (from first to second frame) by VFIformer [Alpher19] and corresponding interpolation. The last two columns: motion field and interpolated frame by enhancing VFIformer with our strategy using RDPs. Our approach results in more satisfactory motion estimation and, thus, better interpolation results.
@@
Restoration-Oriented Video Frame Interpolation (VFI) represents a classic low-level vision task with the objective of augmenting video frame rates by generating intermediary frames that do not exist between consecutive frames. This technique has a wide range of practical applications, such as novel view synthesis [Alpher07], video compression [Alpher08], and cartoon creation [Alpher09]. Nevertheless, frame interpolation continues to present unsolved challenges, including issues related to occlusions, substantial motion, and alterations in lighting conditions. Enhancing the performance of existing restoration-oriented VFI frameworks (not generative models) poses a significant challenge within both the research and industrial communities.
@@
Restoration-oriented VFI methods can be broadly categorized into two main approaches: motion-free [Alpher33, ChannelM2020, Sepcheng2020, Alpher11] and motion-based [jiang2018super, liu2017video, Alpher14, Alpher15, softsplat2020, Alpher17, abme2021, Alpher19, Biformer, fitug, asymmetricble, zhang2023extracting, Vfimamba2024, KIM2025128728], depending on whether they incorporate motion cues like optical flow. Motion-free models typically utilize methods such as kernel prediction or spatial-temporal decoding, which are effective while have limitations, such as being restricted to interpolating frames at fixed time intervals, and their runtime scales linearly with the number of desired output frames. On the other end of the spectrum, motion-based approaches establish dense correspondences between frames and employ warping techniques to generate intermediate pixels. Due to the explicit modeling of temporal correlations, motion-based strategies are more flexible. Moreover, with recent advancements in optical flow technology [Alpher20, Alpher21, Alpher22, Alpher23], motion-based interpolation’s accuracy has been evolved.
Motion estimation between adjacent frames is a pivotal aspect of motion-based VFI. Nevertheless, achieving precise estimation accuracy in existing methods remains a formidable challenge, primarily due to the inherent ambiguity in identifying corresponding areas in adjacent frames for interpolation. This challenge becomes more pronounced when there is a substantial temporal gap in the target video. Previous research has predominantly focused on enhancing estimation accuracy by laboriously evolving network structures. In this paper, we posit that, in addition to network evolution, it is of paramount importance to enhance accuracy by differentiating between various regions prior to the motion estimation process.
In this paper, we present an innovative approach by introducing Region-Distinguishable Priors (RDPs) into motion-based VFI frameworks (restoration-oriented, not generative models that have fidelity concerns). These priors are derived from the existing open-source Segment-Anything Model2 (SAM2) [ravi2024sam2] with minimal impediments. In particular, SAM2 can be applied to videos to segment distinct regions across frames while maintaining consistent representations of these regions over time. The segmentation in each frame enables the estimation of appropriate, region-specific flow values for areas exhibiting different motions, and the temporal consistency of these representations facilitates more accurate motion estimation. These properties satisfies our requirement. Furthermore, we propose a new Hierarchical Region-aware Feature Fusion Module (HRFFM), which is designed to enhance the VFI framework’s encoder, as illustrated in Fig. 2, to refine the corresponding features used in motion estimation. The HRFFM is a plug-and-play module that seamlessly integrates with various motion-based VFI methods without introducing a significant increase in network parameters.
The formulation of RDP from SAM2 is not trivial, as RDP is required to differentiate objects with an arbitrary number, while the output of SAM2 lacks a countable property. To make optimal use of the segment outputs from SAM2 and provide them with the ability to distinguish multiple objects of the same dimensions, we have devised a novel Gaussian embedding strategy for the SAM2 outputs. We employ the SAM2 to produce instance segmentations for two input frames and utilize spatial-varying Gaussian mixtures to transform them into higher-dimensional RDPs. Moreover, SAM2’s temporally consistent segmentation ensures consistent RDP representations across frames. This representation has been demonstrated to outperform naive one-hot encoding or other learnable embedding alternatives.
The obtained temporally-consistent RDPs are integrated into the encoder of the target VFI model, with the primary goal of achieving regional consistency between neighboring frames in VFI. This means that the features of a specific region in two consecutive frames should be similar, which aids in the subsequent motion estimation process. To achieve this objective, HRFFM incorporates RDP into the target model’s hierarchical feature spaces and performs RDP-guided Feature Normalization (RDPFN) in a residual learning fashion to bring target features to desired states. RDPFN is novelly designed to simultaneously harness long- and short-range dependencies to fuse the RDP and image content, enabling the accurate estimation of regional normalization parameters.
@@
Extensive experiments are conducted on public and well-recognized datasets and various VFI networks. It’s verified that our algorithm can bring stable performance improvement consistently on multiple datasets and models. Our strategy produces better motion modeling even with large motion scales, and thus enhances interpolated results (see Fig. 1). In particular, RDPs help produce clearer boundaries and mitigate severe degradations, such as blurring and shape incompleteness. This improvement stems from the fact that our RDPs are derived from SAM2, which provides accurate and well-defined mask boundaries, enabling the complete segmentation of a sufficient number of distinct instances. Meanwhile, the consistency of RDPs between consecutive frames is maintained. In summary, our contribution is three-fold.
  • We underscore the significance of distinguishing different regions within frames to enhance motion estimation and ultimately improve the performance of restoration-oriented VFI. To achieve this, we have innovatively devised a novel and systematic formulation pipeline for RDP using a Gaussian embedding strategy based on the output of SAM2.
  • A new Hierarchical Region-aware Feature Fusion Module is designed to incorporate RDPs into the target model’s encoder. It is a general and plug-and-play strategy for different networks.
@@
Moreover, two key points should be clarified to highlight the role and applicability of our strategy.
The relation between our strategy with existing works. Enhancing interpolation performance can follow two complementary roadmaps: enabling the network to distinguish various regions and evolving the network architecture itself. Updating neural networks alone can be effective, as studied by most existing works [Alpher19, Alpher39, Alpher40, li2023amt, seo2025bim]. However, this approach typically requires large model capacities and extensive training data to handle severe degradations. This is why we propose a new roadmap by incorporating RDPs, which serve as effective priors to further boost the performance of existing VFI networks that have already undergone significant evolution—without introducing substantial additional parameters or requiring large-scale datasets.
The relation between our strategy with generative models. Recently, some approaches have adopted generative models for frame interpolation [wangframer, zhu2025generative, wang2024generative, danier2024ldmvfi]. They primarily rely on pre-trained weights to synthesize frames rather than restore them, focusing on visual quality and aesthetics rather than fidelity. In contrast, our method is designed for restoration-oriented VFI, which emphasizes temporal alignment and leverages RDPs to enhance fidelity. Given the fundamentally different objectives, our approach and generative-model-based strategies belong to distinct categories. Therefore, we have not applied our method to generative models. Additionally, we have not trained or evaluated our method on datasets specifically designed for generative models, such as OpenVid [nan2024openvid]. Moreover, specific tasks like cross-dataset evaluation (e.g., between natural and cartoon videos) are also beyond the scope of our method, as they require strong generative capabilities.
II Related Work
II-A Video Frame Interpolation
@@
Except for diffusion-based models [lyu2025tlbvfitemporalawarelatentbrownian, danier2024ldmvfi, lyu2024fiwcbbd], the current restoration-oriented VFI methods can be broadly categorized into two groups: motion-free and motion-based approaches. Motion-free methods typically create intermediate frames by directly concatenating input frames. Such methods can be further classified into two types: directly-generated methods [ChannelM2020, Alpher25, Alpher26, Alpher27] and kernel-based methods [Sepcheng2020, Alpher29, Alpher30, Alpher31, Alpher32, Alpher33, Alpher34, Alpher35] concerning the generation of intermediate frames. Despite their simplicity, these methods lack robust modeling of motion, making it challenging to align corresponding regions between intermediate frames and input frames. This limitation often results in image blur and the presence of artifacts [Alpher36].
@@
Motion-aware methods explicitly model motion, often represented by optical flow, between two frames to enhance the alignment of distinguishable region information from input frames to intermediate frames. Some early approaches focused solely on predicting inter-frame motion for pixel-level alignment [jiang2018super, liu2017video, liu2019deep]. Subsequent works [Alpher19, Alpher14, softsplat2020, Alpher17, abme2021, Alpher41, sim2021xvfi, reda2022film, zhong2024clearer, liu2023ttvfi, liu2024sparse, liu2022jnmr, hu2023video, hu2024iq, Yoo_2023_ICCV, Biformer, fitug, asymmetricble, zhang2023extracting, Vfimamba2024, KIM2025128728] have introduced separate modules for explicit motion modeling and motion refinement through synthesis, thereby enhancing overall performance. While the current state-of-the-art method has achieved impressive results, these systems still cannot handle practical challenges and need further performance improvement [zhang2023extracting]. Our proposed method offers a novel perspective by incorporating Region Distinguishable Priors into motion-based VFI. Our designed play-and-plug Hierarchical Region-aware Feature Fusion Module provides a straightforward and efficient approach to improving VFI features via RDPs.
II-B Segmentation Models
Traditional segmentation models can be broadly categorized into three types: semantic segmentation, instance segmentation, and panoptic segmentation. The goal of semantic segmentation [zhao2017pyramid, chen2017deeplab] is to assign a predefined category label to every pixel in an image. However, it does not distinguish among different instances of the same class— all objects belonging to the same category are labeled identically. To address this limitation, instance segmentation was introduced [wang2021solo, lee2020centermask], which not only classifies each pixel but also differentiates among individual object instances of the same category. Building upon this, panoptic segmentation [kirillov2019panoptic] was proposed as a unified framework that combines semantic and instance segmentation. It separates image content into “Thing” classes (countable objects, such as people or cars) and “Stuff” classes (amorphous regions, like sky or road). Despite these advances, both instance and panoptic segmentation are limited in scope: not all regions in an image can be meaningfully or distinctly segmented with an instance-aware mask.
@@ II-B Segmentation Models
Recently, the foundational Computer Vision (CV) model for Segment Anything, known as SAM [kirillov2023segment], was recently unveiled. SAM is a substantial Vision Transformer (ViT)-based model that underwent training on an extensive visual corpus (SA-1B). Despite lacking semantic ability, SAM’s capabilities in segmentation have shown promise across various scenarios, underscoring the significant potential of foundational models in the realm of CV. This development marks a groundbreaking stride toward achieving visual artificial general intelligence.
@@
SAM has shown its versatility across several tasks, extending its assistance beyond segmentation. Tasks such as image synthesis [Alpher37] and video super-resolution [Alpher38] have all benefited from SAM’s capabilities. Segment Anything Model 2 (SAM2) [ravi2024sam2] is a foundation model designed for promptable visual segmentation in both images and videos, extending SAM by treating images as single-frame videos. More recently, Segment Anything Model 3 (SAM3) has been introduced as a unified framework that detects, segments, and tracks objects in images and videos based on concept prompts. The continuous evolution of SAM models has greatly advanced the use of segmentation-based priors in real-world videos. In a pioneering effort, we’ve explored SAM2’s potential in VFI, marking the first attempt to apply SAM2 to this domain. Extensive experiments substantiate that SAM2 significantly enhances VFI effectiveness.
In this paper, we use SAM2 to obtain RDP priors instead of other segmentation models, based on a comparative analysis. Semantic segmentation models cannot distinguish between different instances of the same category—for example, two cars moving in opposite directions receive the same label, failing to capture motion differences. Although instance and panoptic segmentation models can separate instances, they are limited by a fixed set of predefined categories and may miss objects not covered by those categories, thus failing to capture all motion patterns. In contrast, SAM2 can segment nearly all objects in an image and distinguish an arbitrary number of instances, making it more suitable for our task. Experiments will further highlight the advantage of using SAM2.
III Method
@@
In this section, we first provide the overview of our strategy in Sec. III-A. Then, two vital components in our framework, i.e., the formulation of Region-Distinguishable Priors and the design of HRFFM, will be elaborated in Sec. III-B and III-C, respectively. One significant component in HRFFM, i.e., RDP-guided Feature Normalization (RDPFN), will be introduced in Sec. III-D. Sec.III-E provides a detailed explanation of the architectures of our networks.
III-A Overview
@@
Implementation. Given $I_{0}$ and $I_{1}$, we first obtain their SAM2 outputs as $M_{0}$ and $M_{1}$. Then, $M_{0}$ and $M_{1}$ (${M_{0},M_{1}}\in\mathbb{R}^{H\times W\times 1}$) are transformed into the desired Region-Distinguishable Priors (RDPs) that can distinguish different regions in frames with a unified representation dimension. Thus, Eq. 1 can be written as
@@
III-B Region-Distinguishable Priors (RDPs)
The drawback of SAM2 outputs for VFI. The original SAM2 model provides segmentation outputs for all instances within an image. SAM2 generates masks for frames, with each pixel value representing an object. Its remarkable segmentation capabilities make it a valuable choice as a region-distinguishable prior. Over time, several variants of SAM2 have been introduced, enhancing its capabilities, including semantic and panoptic segmentation when combined with other models. However, SAM2’s output has limitations when it comes to representing objects with arbitrary numbers, a requirement for RDP. The semantic one-hot embedding is constrained by semantic categories, and the instance one-hot embedding assumes a maximum instance number, making it unable to accommodate new instances during real-world evaluation. Consequently, there is a need to transform SAM2’s output to make it more suitable for RDPs.
Mixture Gaussian embedding strategy. We posit that the representations of segmentation priors can be conceptualized as distributed sampling results with distinct parameters across different regions of an image. These parameters enable the discrimination of regions and the alignment of the same region across multiple frames. In particular, each segmented area can be interpreted as a sampling result from a Gaussian distribution characterized by individual parameters. To facilitate this corresponding sampling process, we begin by establishing a codebook $\mathcal{C}$ that comprises a range of Gaussian parameters, encompassing both mean and variance. Subsequently, each object identified by the SAM2 output can retrieve its specific Gaussian parameters via a hashing mechanism.
We elaborate the procedure and characteristic of the hashing mechanism. First, each region will be assigned an index by SAM2, which can be used to look up the corresponding location in $\mathcal{C}$ with the same index. Then, we can assign unique mean and variance values from $\mathcal{C}$ to each region. In particular, we aim to set a maximum value that accommodates all possible object counts within the frames. This ensures that each region corresponding to an object has a unique entry in the codebook, avoiding collisions with other regions. In this way, not all codebook elements are utilized to generate RDPs; instead, the number of elements used varies across different images, not fixed, depending on the objects present.
In summary, the transformation procedure can be written as: